OpenStudy (anonymous):

Can some one help me? Algebra 2

OpenStudy (anonymous):

Create your own function to teach the Martians about functions. Your function must contain at least two different operations.

OpenStudy (anonymous):

I made 3x + 12

OpenStudy (anonymous):

Using complete sentences, prove to Splott and Fizzle that your function is a legitimate function.

OpenStudy (calculusfunctions):

Your equation is a function as long as the exponent raised to y is one and any term is either positive or negative but not both. eg\[y =\pm \sqrt{x}\]is NOT a function because the square root is preceded by both a positive and a negative sign.

OpenStudy (anonymous):

Thanks @calculusfunctions can you help with a couple more function?

OpenStudy (anonymous):

questions

OpenStudy (calculusfunctions):

Absolutely! What do you need? While you're typing, I'll be back in a couple of minutes. Go ahead.

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Using your function, explain to the Martians how to solve for f(3). Show your work and explain each step using complete sentences.

OpenStudy (anonymous):

F(3) = 3(3) + 12 9 + 12 =21

OpenStudy (anonymous):

Using complete sentences, describe to the Martians how to find the inverse of your function.

OpenStudy (anonymous):

I got x-12/3 Is that right?

OpenStudy (calculusfunctions):

First the above f(3) is correct. Now to find the inverse, i). rewrite y = f(x) (instead of writing f(x), write y), ii). interchange x and y, iii). solve for y, and and iv). rewrite \[y =f ^{-1}(x)\]In other words, change the y into that symbol.

OpenStudy (anonymous):

So instead of 3x + 12, the inverse would be x-12/3 since it's equal to y

OpenStudy (calculusfunctions):

The way you've shown me is incorrect. If you meant (x - 12)/3, then you would be right! \[f ^{-1}(x)=\frac{ x -12 }{ 3 }\]OR\[f ^{-1}(x)=\frac{ 1 }{ 3 }x -4\]

OpenStudy (calculusfunctions):

Understood?!

OpenStudy (anonymous):

Yes, thank you :) one last question

OpenStudy (anonymous):

The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2.

OpenStudy (calculusfunctions):

Yes? Go ahead!

OpenStudy (calculusfunctions):

First, as a rule

OpenStudy (calculusfunctions):

Sorry, I hit post by error. First as a rule \[f(h(x)) \neq h(f(x))\]

OpenStudy (anonymous):

So you're saying the rule is that it will never equal.

OpenStudy (calculusfunctions):

That's correct! However you need to demonstrate that by answering the above question.

OpenStudy (anonymous):

Oh ok, so I will say that x = 4

OpenStudy (anonymous):

h(x) = 5x - 23

OpenStudy (anonymous):

I just make something up, right?

OpenStudy (calculusfunctions):

Boring! Can you not be a tad more creative? Does it always have to be a linear function? But alright go ahead if you must.

OpenStudy (anonymous):

Lol what about h(x) = -6x/ 2 +45

OpenStudy (calculusfunctions):

You do realize that you just gave me another linear function, right?

OpenStudy (anonymous):

No, I truly suck at functions

OpenStudy (calculusfunctions):

LOL. It's okay, I can be patient. After all I'm a teacher. I have to be patient!

OpenStudy (calculusfunctions):

How about a quadratic, cubic, square root, or even a trig or a logarithm function? Anything?

OpenStudy (anonymous):

Um, just to warn you, the more complex the function, stupid is what I become lol. But how about h(x) = 6^2 + x - 42 ?

OpenStudy (anonymous):

Please don't tell me it's still a linear function

OpenStudy (anonymous):

@calculusfunctions ?

OpenStudy (calculusfunctions):

Sorry, I'm experiencing issues with my internet. Please be patient. I'll try to fix it.

OpenStudy (anonymous):

Sorry, I just thought you left

OpenStudy (calculusfunctions):

I know you're joking but never underestimate yourself or sell your self short. If I thought or for one second believed that anyone is stupid as you put it, I wouldn't be here and I wouldn't love teaching as much as I do. You're obviously motivated to learn, which automatically makes you NOT stupid!! You're here to learn and it is my job to teach. Now, you meant to say\[h(x) =6x ^{2}+x -42\]which is a quadratic function. Good finally something different. I thought you just loved linear functions or something. Okay go a head and complete the question and show me your work when you're done.

OpenStudy (calculusfunctions):

Be back in a moment while you're working it out.

OpenStudy (anonymous):

X = 4 h(4) = 6(4)^2 + 4 - 42 h(4) = 6*16 + 4 - 42 = 96 + 4 -42 = 100 - 42 = 58

OpenStudy (anonymous):

Is that how the "6x^2" suppose to have went ?

OpenStudy (calculusfunctions):

Yes that is correct! That's your h(4), now find f(h(4)).

OpenStudy (calculusfunctions):

h(4) = 58 as you said. Thus f(h(4)) = ?

OpenStudy (anonymous):

Ok, here's were I have problems "f(h(4))" I could never truly understand this part, could you explain the steps

OpenStudy (anonymous):

I might have to go soon, thanks for you help :)

OpenStudy (calculusfunctions):

f(h(x)) is what you would call a composition of two functions or better known as composite functions. Composite functions are where the output of the inside function is the input of the outside function. In other words, the outer function is evaluated at the value of the inner function. Hence here your inside function is h(x) and when x = 4, h(x) = 58. Thus f(h(4)) = f(58). Thus now f(58) = ?

OpenStudy (anonymous):

f(58) = h(58)

OpenStudy (calculusfunctions):

Sorry, I'm a slow typer because I like to make sure I'm not making errors. Plus the internet problems don't exactly help either!

OpenStudy (anonymous):

It's ok

OpenStudy (calculusfunctions):

No, that is not what I said. h(4) = 58 so then f(h(4)) = f(58) = 3(58) + 12 = ?

OpenStudy (anonymous):

OH! 58 * 3 = 174 + 12 =186

OpenStudy (calculusfunctions):

YES! Therefore f(h(4)) = 186. Understand?

OpenStudy (anonymous):

Yes, thank you so much :)

OpenStudy (calculusfunctions):

More than welcome. Anytime but you're not finished yet. Now you have to find h(f(4) and show that f(h(4) does not equal h(f(4)).

OpenStudy (anonymous):

:( Ok f(4) = 3(4) + 12 = 12 + 12 =24 h(24) = 6(24)^2 + 4 - 42 6* 576 +4 - 42 3456 +4 -42 3460 -42 = 3418

OpenStudy (calculusfunctions):

Perfect! I hate to say I told you so but see I told you you're smart!

OpenStudy (anonymous):

:)

OpenStudy (calculusfunctions):

Anything else?

OpenStudy (anonymous):

Nope

OpenStudy (calculusfunctions):

Okay then Good Luck and good bye for now. I log in time to time when I have free time from my regular teaching job and I'm bored. You can look for me if you need anything